OpenStudy (anonymous):

One pump can drain a pool in 11 minutes. When a second pump is also used, the pool only takes 9 minutes to drain. How long would it take the second pump to drain the pool if it were the only pump in use?

OpenStudy (anonymous):

Let P = the volume of the pool, a be the rate the first pump drains and b be the second. P/a=11, P/(a+b) = 9 a/P = 1/11, (a+b)/P = 1/9 b/P=1/9-1/11=2/99 P/b=99/2 Therefore, pump b drains the pool in 99/2 minutes. I think

OpenStudy (anonymous):

w/a=11 w/(a+b)=9 w/b=? a=w/11 w/9=a+b, w/9=w/11+b w/9-w/11=b, 2w/99=b w/(2w/99)=t(b) t(b)=99/2=49.5

OpenStudy (anonymous):

Moory is correct. However,assuming x be the required answer,it's could be simply be solved like this : \[ -\frac1{11} - \frac1x = -\frac 19\Rightarrow -\frac1x =-\frac2{99}\Rightarrow x = 49.5\]